OpenStudy (anonymous):

Need some help! Medals to anyone who answers! Thanks! It is much appreciated! :D Part 1: Decide whether you would use the graphing, substitution, or elimination method to solve the following system of equations. Explain, in complete sentences, why you chose that method. Part 2: Solve the following system of equations and show all of your work. 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

I would suggest using elimination because you have a 5y in the first equation and a -5y in the second equation. You can add the two equations together to eliminate the y variable and solve for x. Once you solve for x, then you can substitute that value into one of the original equations to get y.

OpenStudy (anonymous):

2x + 5y = -3 -15x - 5y = -10 2x + (-15x) + 5y + (-5y) = -3 + (-10) -13x = -13 x = 1 2(1) + 5y = -3 5y = -5 y = -1